### v4

@@ C.2   Repeated cross sections
where $\delta_{i,e,l}^{r}=Y_{i}-\theta_{e,l}^{r}\cdot D_{i}$.
C.3   Triple DID-IV designs
We also present the influence function of our triple DID-IV estimator when panel or repeated cross section data are available. Let $\theta_{e,l,g}^{p}$ and $\theta_{e,l,g}^{r}$ define
$\theta_{e,l,g}^{p}\equiv\frac{\alpha_{e,A}-\alpha_{e,B}-[\beta_{U,A}-\beta_{U,B}]}{\pi_{e,A}-\pi_{e,B}-[\gamma_{U,A}-\gamma_{U,B}]},$
$\theta_{e,l,g}^{r}\equiv\frac{[(\alpha_{e,e+l,A}-\alpha_{e,e-1,A})-(\alpha_{e,e+l,B}-\alpha_{e,e-l,B})]-[(\beta_{U,e+l,A}-\beta_{U,e-1,A})-(\beta_{U,e+l,B}-\beta_{U,e-1,B})]}{[(\pi_{e,e+l,A}-\pi_{e,e-1,A})-(\pi_{e,e+l,B}-\pi_{e,e-l,B})]-[(\gamma_{U,e+l,A}-\gamma_{U,e-1,A})-(\gamma_{U,e+l,B}-\gamma_{U,e-1,B})]},$
where
$\alpha_{e,g^{\prime}}=E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e,g_{i}=g^{\prime}],\hskip 5.69054pt\beta_{U,g^{\prime}}=E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U,g_{i}=g^{\prime}],$
$\pi_{e,g^{\prime}}=E[D_{i,e+l}-D_{i,e-1}|E_{i}=e,g_{i}=g^{\prime}],\hskip 5.69054pt\gamma_{U,g^{\prime}}=E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U,g_{i}=g^{\prime}]\hskip 5.69054pt(g^{\prime}=A,B),$
and
$\alpha_{e,t,g^{\prime}}=E[Y_{i}|E_{i}=e,T_{i}=t,g_{i}=g^{\prime}],\hskip 5.69054pt\beta_{U,t,g^{\prime}}=E[Y_{i}|E_{i}\in U,T_{i}=t,g_{i}=g^{\prime}],$
$\pi_{e,t,g^{\prime}}=E[D_{i}|E_{i}=e,T_{i}=t,g_{i}=g^{\prime}],\hskip 5.69054pt\gamma_{U,t,g^{\prime}}=E[D_{i}|E_{i}\in U,T_{i}=t,g_{i}=g^{\prime}]\hskip 5.69054pt(g^{\prime}=A,B,\hskip 5.69054ptt=e+l,e-1).$
In panel data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{p}$ and has the following influence function.
$\psi_{i,e,l,g}^{p}=\frac{1}{DID_{e,l,g}^{p}}(\zeta_{i,e,A}-\zeta_{i,e,B}-\zeta_{i,U,A}+\zeta_{i,U,B}),$
where $DID_{e,l,g}^{p}=\pi_{e,A}-\pi_{e,B}-[\gamma_{U,A}-\gamma_{U,B}]$, $\delta_{i,e,l,g}^{p}=(Y_{i,e+l}-Y_{i,e-1})-\theta_{e,l,g}^{p}\cdot(D_{i,e+l}-D_{i,e-1})$, and
$\zeta_{i,e,g^{\prime}}=\frac{\mathbf{1}_{e,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{p}-E[\delta_{i,e,l,g}^{p}|E_{i}=e,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{e,g^{\prime}}]},\hskip 5.69054pt\zeta_{i,U,g^{\prime}}=\frac{\mathbf{1}_{U,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{p}-E[\delta_{i,e,l,g}^{p}|E_{i}\in U,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{U,g^{\prime}}]}.$
In repeated cross section data settings, our triple DID-IV estimator is the sample analog of $\theta_{e,l,g}^{r}$ and has the following influence function.
$\psi_{i,e,l,g}^{r}$
$=\frac{1}{DID_{e,l,g}^{r}}([\zeta_{i,e,e+l,A}-\zeta_{i,e,e-1,A}-\zeta_{i,e,e+l,B}+\zeta_{i,e,e-1,B}]-[\zeta_{i,U,e+l,A}-\zeta_{i,U,e-1,A}-\zeta_{i,U,e+l,B}+\zeta_{i,U,e-1,B}]),$
where $DID_{e,l,g}^{r}=[(\pi_{e,e+l,A}-\pi_{e,e-1,A})-(\pi_{e,e+l,B}-\pi_{e,e-l,B})]-[(\gamma_{U,e+l,A}-\gamma_{U,e-1,A})-(\gamma_{U,e+l,B}-\gamma_{U,e-1,B})]$, $\delta_{i,e,l,g}^{r}=Y_{i}-\theta_{e,l,g}^{r}\cdot D_{i}$, and
$\zeta_{i,e,t,g^{\prime}}=\frac{\mathbf{1}_{e,t,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{r}-E[\delta_{i,e,l,g}^{r}|E_{i}=e,T_{i}=t,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{e,t,g^{\prime}}]},$
$\zeta_{i,U,t,g^{\prime}}=\frac{\mathbf{1}_{U,t,g^{\prime}}\cdot\left[\delta_{i,e,l,g}^{r}-E[\delta_{i,e,l,g}^{r}|E_{i}\in U,T_{i}=t,g_{i}=g^{\prime}]\right]}{E[\mathbf{1}_{U,t,g^{\prime}}]}.$
Appendix D Decomposing the TWFEIV estimand in Oreopoulos (2006)
Here, we quantify the bias terms in the TWFEIV estimand, arising from the bad comparisons in Oreopoulos (2006). Let $e=1947$ and $U=1957$ denote England and North Ireland respectively, and let $N_{e,t}$ denotes the sample size for England in cohort $t$. Let $R_{e,t}$ denote the sample mean of random variable $R_{i,t}$ in region $e$ ($e=1947,1957$) and cohort $t$:
$R_{e,t}\equiv\frac{1}{N_{e,t}}\sum_{i}^{N_{e,t}}R_{e(i),t}.$
First, we decompose the TWFEIV estimator $\hat{\beta}_{IV}$ as follows [^11] For the detailed calculation steps, see the proof of Lemma 7 in Miyaji (2024). In our decomposition, the reference period ($t=1$) in the proof is $t=1946$.:
@@ Appendix D Decomposing the TWFEIV estimand in Oreopoulos (2006)
Using similar arguments in the proof of Theorem 4, one can show that in staggered DID-IV designs, we have
@@
where $CAET^{1}_{e,t}=E[D_{i}-D_{i}^{\infty}|T_{i}=t,E_{i}=e]$. Similarly, we have
@@
Combining the result (30) with (34) and by Slutsky’s theorem, we obtain the following decomposition result for the TWFEIV estimand in Oreopoulos (2006):
where $\Delta_{e,t}$ is:
$\displaystyle\frac{CAET^{1}_{e,t}\cdot CLATT_{e,t}-CAET^{1}_{U,t}\cdot CLATT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
and the weights $w^{1}_{e,t}$ and $w^{2}_{e,t}$ are:
$w_{e,t}^{1}=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (36)
$w_{e,t}^{2}=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CATT^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (37)
In Oreopoulos (2006), we can only identify $CLATT_{e,t}$ between $1947$ and $1956$ because the policy change occurred in England in $1947$ and in North Ireland in $1957$. This implies that each $\Delta_{e,t}$ in equation (35) is the bias term arising from the bad comparisons between $1957$ and $1965$ performed by the TWFEIV regression.
Figure 4 plots the weight and the corresponding estimate for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term). The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CLATT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, using equations (30) - (34).
Figure 4 shows that all the CLATTs are positive and positively weighted, whereas all the bias terms are positive and negatively weighted. This indicates that the TWFEIV estimand is negatively biased in the setting of Oreopoulos (2006).
@@ Appendix E Comparing DID-IV to Fuzzy DID
In this appendix we compare DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018) (henceforth, ”dCDH”). We assume that the reader has briefly read dCDH. We first explain how dCDH’s identifying assumptions are related to ours in this article. We then point out the issues inherent in Fuzzy DID designs. First, we show that dCDH’s assumptions ex ante impose strong and asymmetric restrictions on treatment adoption behavior across units between exposed and unexposed groups. Next, under these restrictions, we show that dCDH’s target parameter, the SLATET, can be decomposed into a weighted average of two different causal parameters. At the end of this appendix, we also revisit the ”issue” pointed out by dCDH regarding the use of the Wald-DID estimand, and show that while their argument is correct, it is misguided.
@@ E.1   Comparing the identifying assumptions in dCDH and this article
In this section, we investigate the detailed connections between dCDH’s identifying assumptions and ours. Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we refer to Assumptions 14-18. For now, we ignore the difference in the definition of target parameter between dCDH and this article.
dCDH consider the following identification assumptions for the Wald-DID estimand to capture their target parameter, the SLATET [^13] In repeated cross section settings, the formal definition is provided in Assumption 29.. In dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation [^14] This follows from two observations in dCDH. First, dCDH call the groups $G=1$ and $G=0$ treatment and control groups, respectively. Second, the group variable $G$ is included in Assumption 26 (Treatment participation equation).. We therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively. [^15] Note that in Appendix B.2, we construct the group variable $E$ from the instrument path $Z$ to uncover Assumption 16 (No anticipation in the first stage), the hidden assumption in the previous literature. For the detailed explanations for each assumption, see dCDH.
@@
We now state the relationships between dCDH’s assumptions and ours. We first explain the similarities. First, both dCDH and this article impose the condition (39) in Assumption 24 (Fuzzy design), as Assumptions 15-17 imply this condition [^16] From Assumptions 15-17, we have: $E[D|E=1,T=1]-[D|E=1,T=0]-(E[D|E=0,T=1]-E[D|E=0,T=0])$ $=$ $E[D_{1}((0,1))-D_{1}((0,0))|E=1,T=1]$ $=$ $Pr(CM^{Z}|E=1,T=1)>0.$
Next, the treatment participation equation in Assumption 26 corresponds to Assumption 15 (Monotonicity) in this article. The equivalence between the treatment participation equation and the potential treatment choices framework with Assumption 15 (Monotonicity) follows from the result of Vytlacil (2002). We point out that dCDH implicitly assume the sharp assignment of the instrument $Z$ in their proofs: they assume $v_{10}=v_{00}$ in Assumption 26; that is, the thresholds in the treatment participation equation are the same between exposed and unexposed groups in period 00.
Finally, dCDH implicitly impose Assumption 14 (Exclusion restriction) and Assumption 16 (No anticipation in the first stage) when they formalize potential outcomes and treatment participation equations. Specifically, dCDH do not introduce the instrument path in their framework and define potential outcomes as $Y(D)$, which implies Assumption 14. As one can see in Assumption 26, dCDH assume that the treatment participation equation only depends on the current instrument status, which implies Assumption 16 since dCDH assume the sharp assignment of the instrument: $v_{10}=v_{00}$.
The differences between dCDH’s assumptions and ours are as follows. First, dCDH assume the parallel trends assumption in the untreated outcome (Assumption 27), whereas we assume the parallel trends assumption in the unexposed outcome. Second, dCDH assume the stable treatment rate assumption in an unexposed group (Assumption 25), whereas we assume the parallel trends assumption in the treatment between exposed and unexposed groups. Third, dCDH impose the independence assumption between unobserved heterogeneity $V$ and time variable $T$ given $Z$ in Assumption 26, whereas we do not impose such restriction. Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 28) to capture the SLATET, whereas we do not require such restriction for this estimand to identify the LATET. Finally, dCDH assume the condition (38) in Assumption 24 (Fuzzy design), which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
@@
where we define171717dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively. When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
@@
This assumption requires that the treatment effects among the switchers should be the same between exposed and unexposed groups, which dCDH call the ”homogeneous” treatment effect assumption. In Appendix E.4, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the ”homogeneous” treatment effect between the two groups. In the following discussion, we therefore treat Assumptions 24-28 as dCDH’s main identifying assumptions.
E.2   Strong and asymmetric restrictions in Fuzzy DID designs
We show that dCDH’s identifying assumptions implicitly impose strong and asymmetric restrictions on treatment adoption behavior across units between exposed and unexposed groups. First, we prove some lemmas implied by dCDH’s assumptions. Based on these lemmas, we then create the tables that present all the types in exposed and unexposed groups under Fuzzy DID designs. These tables show that dCDH’s assumptions ex ante exclude many types in both groups and these restrictions are heavily assigned on an unexposed group. At the end, we provide the empirical implication of the distinction between dCDH’s restrictions and ours in DID-IV designs.
First, we show that Assumption 26 (Treatment participation equation) implies the following assumption.
@@ E.2   Strong and asymmetric restrictions in Fuzzy DID designs
because we assume that the unobserved heterogeneity $V$ is independent of time variable $T$ conditional on the instrument $Z$.
Completing the proof. ∎
Lemma 1 shows that in Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively. Note that the treatment participation equation itself does not imply Assumption 30: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption. Indeed, in this article, we do not assume this restriction.
As we have seen in section 2.4 in this article, time also plays the role of instrument in DID-IV settings. From this observation, we can view that in an unexposed group ($z=0$), Assumption 30 corresponds to a monotonicity assumption with respect to time, formally defined below.
@@
We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Tables 9 - 10 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group [^18] In both tables, we can only observe either $D_{0}(z)$ or $D_{1}(z)$ in group $z$ as we consider repeated cross section settings.: five and three types exist in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
The distinction between dCDH’s restrictions and ours is apparent when we compare Tables 9 - 10 to Table 1 - 2191919Tables 1-2 are constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings. The only modification is that as in Tables 9-10, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
Table 9: Exposed group ($z=1$)
observed counterfactual
$D_{0}(0)$  or  $D_{1}(1)$ $D_{1}(0)=1$ $D_{1}(0)=0$
$D_{0}(0)=1,D_{1}(1)=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}(0)=1,D_{1}(1)=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(1)=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}(0)=0,D_{1}(1)=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 10: Unexposed group ($z=0$)
observed counterfactual
$D_{0}(0)$  or  $D_{1}(0)$ $D_{1}(1)=1$ $D_{1}(1)=0$
$D_{0}(0)=1,D_{1}(0)=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}(0)=1,D_{1}(0)=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}(0)=0,D_{1}(0)=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}(0)=0,D_{1}(0)=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
These differences between dCDH’s restrictions and ours seem critical in many empirical settings. Unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions seem difficult to assess from the institutional knowledge in a given application, and the practitioners may want to avoid imposing these restrictions in empirical work. For instance, in some cases, we may have no valid reason for excluding the specific type $NT^{Z}\land DF^{T}$ in an exposed group, who are not affected by instrument but affected by time. Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.
Although these restrictions in dCDH seem strong, one might argue that each assumption, such as the condition (38) in Assumption 24 and Assumption 25 (Stable percentage of treated units in an unexposed group), can be justifiable in certain applications. However, these assumptions give rise to two additional issues. First, when we impose the condition (38) in Assumption 24, we should test its plausibility from the data because this condition excludes the case that the treatment rate in an exposed group is stable during the two periods: $E[D|Z=1,T=1]=E[D|Z=1,T=0]$. Although dCDH argue that ”Assumption $1$ (Assumption 24 in this article) is just a way to define the treatment and the control group (exposed and unexposed group) in our fuzzy setting”, they do not explain how we should proceed when this condition is violated in practice.
Second, when we impose Assumption 25, we should find or estimate the groups whose treatment rates are stable over time. Although dCDH propose the two-step estimation method when we should estimate the stable treatment group from the data, its performance and feasibility are not clear, especially for multiple period and multiple group settings.

### v5

@@ Appendix B Deriving the influence function in repeated cross section settings
where $\mathbf{1}_{A}$ is the indicator function and takes one if $A$ is true, and $\delta_{i,e,l}^{r}=Y_{i}-\theta_{e,l}^{r}\cdot D_{i}$.
Appendix C Decomposing the TWFEIV estimand in Oreopoulos (2006)
[…]
@@ Appendix C Decomposing the TWFEIV estimand in Oreopoulos (2006)
where we define:
$CAET^{1}_{e,t}=E[D_{i}-D_{i}^{\infty}|T_{i}=t,E_{i}=e],$
$\Delta_{e,t}=\displaystyle\frac{CAET^{1}_{e,t}\cdot CLATT_{e,t}-CAET^{1}_{U,t}\cdot CLATT_{U,t}}{CAET^{1}_{e,t}-CAET^{1}_{U,t}},$
and the weights $w^{1}_{e,t}$ and $w^{2}_{e,t}$ are:
$w_{e,t}^{1}$ (21)
$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (22)
$w_{e,t}^{2}$ (23)
$=\frac{E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CATT^{1}_{U,t})}{\sum\limits_{U-1\geq t\geq e}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot CAET^{1}_{e,t}+\sum\limits_{t\geq U}E[\hat{Z}_{i,t}|E_{i}=e]\cdot n_{e,t}\cdot(CAET^{1}_{e,t}-CAET^{1}_{U,t})},$ (24)
@@
Figure 3 plots the weight and the corresponding estimate for each $CLATT_{e,t}$ and $\Delta_{e,t}$ (bias term). The consistent estimators for each $w^{1}_{e,t}$, $w^{2}_{e,t}$, $CLATT_{e,t}$ and $\Delta_{e,t}$ are constructed from the sample analogue, exploiting $\widehat{WDID}_{e,t}$ and $(D_{e,t}-D_{e,1946})-(D_{U,t}-D_{U,1946})$.
[…]
@@
Appendix D Relationship between DID and DID-IV
In this appendix, we describe the relationship between DID and DID-IV designs, and establish that DID-IV is DID with noncompliance of the assigned treatment.
$D_{i,0}((0,1))=D_{i,0}((0,0))=0\hskip 8.53581pt\text{for all $i$}.$ (25)
These restrictions imply that we have Tables 5-6 under DID designs: there exists only the type $CM^{Z}\land NT^{T}$ in exposed and unexposed groups, respectively.
Second, the target parameter in DID designs, the ATT is equal to the LATET in the canonical DID set up; this follows from Tables 5-6:
Table 5: Exposed group (assigned to $z=(0,1)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=0$
$D_{0}((0,0))=0,D_{1}((0,1))=1$ $CM^{Z}\land NT^{T}$
Table 6: Treatment group (assigned to $z=(0,0)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$
$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$
$ATT$ $=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1]$
$=E[Y_{i,1}(1)-Y_{i,1}(0)|E_{i}=1,CM^{Z}]$
$=LATET.$
Finally, a well-known parallel trends assumption in DID designs can also be interpreted as requiring the same expected time gain between exposed and unexposed groups. From Tables 5-6, we can rewrite the parallel trends assumption in the untreated outcome as:
$E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=0]=E[Y_{i,1}(0)-Y_{i,0}(0)|E_{i}=1]$
$\iff$ $\Delta_{NT^{T},0}=\Delta_{NT^{T},1}$
$\iff$ $\Delta_{0}=\Delta_{1},$
[…]
@@ Appendix E Comparing DID-IV to Fuzzy DID
[…]
@@ E.1   Comparing the identifying assumptions in dCDH and this article
In this subsection, we investigate the detailed connections between dCDH’s identifying assumptions and ours. Because dCDH consider the repeated cross section settings, when we compare the identifying assumptions between the two papers, we refer to Assumptions 14-19. For now, we ignore the difference in the definition of target parameter.
dCDH consider the following identification assumptions for the Wald-DID estimand to capture the SLATET. [^25] In repeated cross section settings, the formal definition is provided in Assumption 31. Note that in dCDH, a group variable $G$ plays the role of instrument $Z$ in our notation. [^26] This follows from two observations in dCDH. First, dCDH call the groups $G=1$ and $G=0$ treatment and control groups, respectively. Second, the group variable $G$ is included in Assumption 28 (Treatment participation equation). We therefore replace $G$ with $Z$ and call the groups $Z=1$ and $Z=0$ exposed and unexposed groups, respectively. [^27] Note that in Appendix A.2, we distinguish the group variable $E$ from the instrument path $Z$ to uncover Assumption 16 (No anticipation in the first stage). For the detailed explanations for each assumption, see dCDH.
@@
[…]
Finally, dCDH implicitly impose Assumption 14 and Assumption 16 when they formalize potential outcomes and treatment participation equations. Specifically, dCDH do not introduce the instrument path in their framework and define potential outcomes as $Y(D)$, which implies Assumption 14. As one can see in Assumption 28, dCDH assume that the treatment participation equation only depends on the current instrument status, which implies Assumption 16 since dCDH assume the sharp assignment of the instrument: $v_{10}=v_{00}$.
The differences between dCDH’s assumptions and ours are as follows. First, dCDH assume the parallel trends assumption in the untreated outcome (Assumption 29), whereas we assume the parallel trends assumption in the unexposed outcome. Second, dCDH assume the stable treatment rate assumption in an unexposed group (Assumption 27), whereas we assume the parallel trends assumption in the treatment between exposed and unexposed groups. Third, dCDH impose the independence assumption between unobserved heterogeneity $V$ and time variable $T$ given $Z$ in Assumption 28, whereas we do not impose such restriction. Forth, dCDH argue that the Wald-DID estimand requires the stable treatment effect assumption during the two periods (Assumption 30) to capture the SLATET, whereas we do not require such restriction for this estimand to identify the LATET. Finally, dCDH assume the condition (27) in Assumption 26, which requires that the treatment rate in an exposed group increases between the two periods, whereas we do not assume this condition.
@@
where we define282828dCDH denote SLATET and SLATET’ as $\Delta$ and $\Delta^{\prime}$ respectively. When we define the SLATET, we use $D_{0}(0)=D_{0}(1)$, which follows from the sharp assignment of the instrument: $v_{10}=v_{00}$.:
@@
This assumption requires that the treatment effects among the switchers should be the same between exposed and unexposed groups, which dCDH call the homogeneous treatment effect assumption. In Appendix E.4, we present the decomposition result for the SLATET, and show that we cannot straightforwardly interpret this assumption as requiring the homogeneous treatment effect between the two groups. In the following discussion, we therefore treat Assumptions 26-30 as dCDH’s main identifying assumptions.
E.2   Strong restrictions under Fuzzy DID designs
In this subsection, we show that dCDH’s identifying assumptions ex ante impose strong restrictions on treatment adoption behavior across units. First, we prove some lemmas implied by dCDH’s assumptions. Based on these lemmas, we then create the tables that present all the types in exposed and unexposed groups under Fuzzy DID designs. These tables show that dCDH’s assumptions ex ante exclude many types in both groups and these restrictions are heavily assigned on an unexposed group. At the end, we provide the empirical implication of the distinction between dCDH’s restrictions and ours.
First, we show that Assumption 28 implies the following assumption.
@@ E.2   Strong restrictions under Fuzzy DID designs
[…]
because we assume that the unobserved heterogeneity $V$ is independent of time variable $T$ conditional on the instrument $Z$. ∎
Lemma 1 shows that under Fuzzy DID designs, the treatment status changes between period 00 and period $1$ in a monotone (uniform) way in exposed and unexposed groups, respectively. Note that the treatment participation equation itself does not imply Assumption 32: the independence between the unobserved heterogeneity $V$ and time variable $T$ given $Z$ yields this assumption. Indeed, in this article, we do not assume this restriction.
As we have seen in Subsection 2.4, time also plays the role of instrument in DID-IV settings. From this observation, we can view that in an unexposed group ($z=0$), Assumption 32 corresponds to a monotonicity assumption with respect to time, formally defined below.
@@
We now clarify how dCDH’s identifying assumptions restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Tables 3-4 in Subsection 2.5 show that their identifying assumptions exclude many types in both groups and the restrictions are heavily imposed on an unexposed group: five and three types exist in exposed and unexposed groups respectively, where the types painted in gray color are excluded in both groups.
Comparing Tables 3-4 to Tables 1-2, one can see the difference between dCDH’s restrictions and ours [^29] Tables 1-2 are constructed under panel data settings, but the types represented in these tables remain consistent even when considering repeated cross section settings. The only modification is that as in Tables 3-4, one can only observe either $D_{0}(z)$ or $D_{1}(z)$ for unit $i$ in group $z$ in each table.: in both groups, we only exclude the defiers $DF^{Z}$ by Assumption 15 (Monotonicity), whereas dCDH additionally exclude the type $NT^{Z}\land DF^{T}$ in an exposed group by Lemma 3, and additionally exclude the time defiers $DF^{T}$ and the time compliers $CM^{T}$ in an unexposed group by Lemma 2.
The distinction between dCDH’s restrictions and ours appears to be important in many empirical settings. Unlike the monotonicity assumption, which eliminate the compliers $CM^{Z}$ in both groups, dCDH’s additional restrictions might be difficult to defend in some applications, and the practitioners may want to avoid imposing these restrictions in practice. For instance, in some cases, we may have no valid reason for excluding the specific type $NT^{Z}\land DF^{T}$ in an exposed group, who are not affected by instrument but affected by time. Similarly, in other cases, we may have no justification to exclude the time defiers $DF^{T}$ and time compliers $CM^{T}$ in an unexposed group because of the steady growth of the treatment rate.